OpenStudy (anonymous):

Find the Coordinates of the points of intersection for the graphs of x^2+2y^2=33 and x^2+y^2=2+19

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

set them equal, solve for \(x\)

OpenStudy (misty1212):

\[x^2+2y^2=33\\ x^2+y^2=21\] like that?

OpenStudy (misty1212):

the \(2+19\) is a little confusing is that really what it says?

OpenStudy (anonymous):

2x+19 sorry!!!

OpenStudy (misty1212):

oh that sucks we can still do it

OpenStudy (misty1212):

\[x^2+2y^2=33\\ x^2+y^2=2x+19\]

OpenStudy (misty1212):

i don't know of a quick way to do it if you subtract the second equation from the first you get \[y^2=-2x+14\]

OpenStudy (misty1212):

then you can substitute that in to either equation to get an equation with only \(x\)

OpenStudy (anonymous):

I am doing test corrections so I already know the answer needs to be (5, +-2) (-1,4). I just need help on the work. If having the answer helps at all

OpenStudy (misty1212):

for example here \[x^2+2y^2=33\\x^2+2(-2x+14)=33\]

OpenStudy (misty1212):

yeah i know the answer, used wolfram already just trying to figure out how to do it is all in any case once you have \[x^2+2(-2x+14)=33\] it is a quadratic equation with only \(x\) and you will get two x values do you know how to solve it?

OpenStudy (anonymous):

no clue

OpenStudy (misty1212):

lol multiply out first on the left \[x^2-4x-28=33\]

OpenStudy (misty1212):

oops \[x^2-4x+28=33\] thats better

OpenStudy (misty1212):

subtract \(33\) from both sides get \[x^2-4x-5=0\] then factor

OpenStudy (anonymous):

im trying to get x right?

OpenStudy (misty1212):

yes, you have to factor it first you know what i mean?

OpenStudy (anonymous):

like get to 4x=x^2-5 ?

OpenStudy (misty1212):

oh no

OpenStudy (misty1212):

like \[x^2-x-6=(x-3)(x+2)\] like that factoring

OpenStudy (misty1212):

you know how to factor?

OpenStudy (anonymous):

It has been awhile.

OpenStudy (misty1212):

think of two numbers that multiply to get \(-5\) and add to get \(-4\) hint, you have almost no choices

OpenStudy (anonymous):

5 and -1

OpenStudy (misty1212):

close they multiply to get \(-5\) but if you add \(5+(-1)\)you get \(4\) and you want \(-4\)

OpenStudy (misty1212):

just change the signs

OpenStudy (anonymous):

-5 and 1

OpenStudy (misty1212):

right, so \[x^2-4x-5=(x-4)(x+1)\]

OpenStudy (misty1212):

solve \[(x-5)(x+1)=0\] by solving \[x-5=0\] and \[x+1=0\] those will give the two answers you knew you were supposed to get

OpenStudy (anonymous):

so do i do the same thing with y now